I'm looking for some needed assistance to help solve this equation by the addition method. Can anyone help? Thank you.
3x -2.5y = 7.125
2.5x - 3y = 7.3125

Answer by mananth(16949)

3x -2.5y = 7.125..
2.5x - 3y = 7.3125
..
Multiply equation 1 by -2.5 and second eauation by 3
..
-7.5x+6.25y= -17.8125
7.5x -9y = 21.9375
..
add them x cancels off
-2.75y = 4.125
y= -1.5
plug the value of y in first equation
3*(x)-2.5(-1.5)=7.125
3x= 7.125+3.75= 7.125
x= 1.125